OpenStudy (anonymous):

How do you implicitly differentiate: sqrt(xy)=x^2y+1

OpenStudy (amistre64):

a derivative is a derivative is a derivative; the rules of differentiation are the same no matter what your looking at

OpenStudy (anonymous):

Is the answer: \[y'=(2x)/(2\sqrt(xy))\]

OpenStudy (amistre64):

dunno, what steps did you take to come to that conclusion?

OpenStudy (anonymous):

\[1/2(xy)^{-1/2}=2xy'\] then \[1/2\sqrt(xy)=2xy'\]

OpenStudy (anonymous):

I used only product and power rule.

OpenStudy (amistre64):

sqrt(xy) = x^2 y + 1 sqrt(xy)' = (x^2 y)' + 1' sqrt(xy)' (xy)'= (x^2 y)' + 1' (xy)'/2sqrt(xy) = x'^2 y + x^2 y' (x'y+xy')/2sqrt(xy) = 2x x' y + x^2 y' if we assume that this is diffed wrt x then x'=1 (y+xy')/2sqrt(xy) = 2xy + x^2 y'

OpenStudy (amistre64):

(y+xy')/2sqrt(xy) = 2xy + x^2 y' (y+xy')/2sqrt(xy) - x^2 y'= 2xy y' (y+x)/2sqrt(xy) - x^2= 2xy y' = 2xy/((y+x)/2sqrt(xy) - x^2) \[ \large y' = \cfrac{2xy}{\frac{y+x}{2\sqrt{xy}} - x^2}\]